QUESTION 4. Let A and B be the following vectors, A = 4N at 60º and B = 3.5 N at 150º, where “N” stands for Newtons. Determine:

A) A + B B) A – B C) 4A + 3B.

Please I need a right explanation, step by step.
Answer #1

here,

A = 4 N * ( cos(60) i + sin(60) j)

A = 2 N i + 3.46 N j

B = 3.5 N * ( cos(150) i + sin(150) j)

B = - 3.03 N i + 1.75 N j

a)

R = A + B

R = - 1.03 N i + 5.21 N j

the magnitude of R , |R| = sqrt(1.03^2 + 5.21^2) = 5.31 N

direction , theta = arctan(5.21 /( - 1.03))

theta = 101.18 degree counterclockwise from +X axis

b)

R = A - B

R = 5.03 N i + 1.71 N j

the magnitude of R , |R| = sqrt(5.03^2 + 1.71^2) = 5.31 N

direction , theta = arctan(1.71 /(5.03))

theta = 18.8 degree counterclockwise from +X axis

c)

R = 4A + 3B

R = - 1.09 N i + 19.09 N j

the magnitude of R , |R| = sqrt(1.09^2 + 19.09^2) = 19.2 N

direction , theta = arctan(19.09 /( -1.09))

theta = 93.3 degree counterclockwise from +X axis